# Connecting Federated ADMM to Bayes

### v1

@@ Appendix F Additional experimental results
Table 2: Mean accuracy (standard deviation in parentheses) over three runs (except CIFAR-10), after a fixed number of communication rounds. For each run, we use the maximum accuracy over the previous 3 rounds (mean accuracy over previous 3 rounds in Table 1). We bold the top two performing algorithms (even if their standard deviations overlap with others). We see that FedLap performs at least as good as FedDyn in all settings (with a bigger difference when the global model performs better with a non-zero weight decay, like in UCI Credit). FedLap-Cov significantly improves upon FedLap in most settings (except for MNIST homogoeneous, where all methods perform equally well), and FedLap-Func also improves performance (especially seen on CIFAR10).
@@
Dataset Comm FedAvg FedProx FedDyn FedLap FedLap FedLap
Round -Cov -Func
@@
Table 4: Mean accuracy (standard deviation in parentheses) over three runs, after a fixed number of communication rounds, on a homogoneous split of UCI Credit into $K=10$ clients. For each run, we use the average accuracy over the previous 3 rounds to account for instabilities (maximum accuracy over previous 3 rounds in Table 2), and bold the top two performing algorithms (even if their standard deviations overlap with others). We see that FedLap and FedLap-Cov perform the best.

### v2

@@ Appendix F Additional experimental results
In Table 5 we report average test negative log-likelihood results (instead of accuracy) at the end of training. We see that on average, FedLap performs at least as well as FedDyn (sometimes significantly outperforming FedDyn, such as on FMNIST (homog), FMNIST (heterog) and CIFAR10). FedLap-Cov and FedLap-Func improve even further, as might be expected due to the Bayesian interpretation.
@@
Table 2: Mean accuracy (standard deviation in parentheses) over three runs (except CIFAR-10), after a fixed number of communication rounds. For each run, we use the maximum accuracy over the previous 3 rounds (mean accuracy over previous 3 rounds in Table 1). We bold the top two performing algorithms (even if their standard deviations overlap with others). We see that FedLap performs at least as good as FedDyn in all settings (with a bigger difference when the global model performs better with a non-zero weight decay, like in UCI Credit). FedLap-Cov significantly improves upon FedLap in most settings (except for homogeneous MNIST, where all methods perform equally well), and FedLap-Func also improves performance (especially seen on CIFAR10).
@@
Dataset Comm FedAvg FedProx FedDyn FedLap FedLap
Round -Cov
@@
Table 4: Mean accuracy (standard deviation in parentheses) over three runs, after a fixed number of communication rounds, on a homogoneous split of UCI Credit into $K=10$ clients. For each run, we use the average accuracy over the previous 3 rounds to account for instabilities, and bold the top two performing algorithms (even if their standard deviations overlap with others). We see that FedLap and FedLap-Cov perform the best.